### v1

# Tilt your Head: Activating the Hidden Spatial-Invariance of Classifiers

Abstract
Deep neural networks are applied in more and more areas of everyday life. However, they still lack essential abilities, such as robustly dealing with spatially transformed input signals. Approaches to mitigate this severe robustness issue are limited to two pathways: Either models are implicitly regularised by increased sample variability (data augmentation) or explicitly constrained by hard-coded inductive biases. The limiting factor of the former is the size of the data space, which renders sufficient sample coverage intractable. The latter is limited by the engineering effort required to develop such inductive biases for every possible scenario. Instead, we take inspiration from human behaviour, where percepts are modified by mental or physical actions during inference. We propose a novel technique to emulate such an inference process for neural nets. This is achieved by traversing a sparsified inverse transformation tree during inference using parallel energy-based evaluations. Our proposed inference algorithm, called Inverse Transformation Search (ITS), is model-agnostic and equips the model with zero-shot pseudo-invariance to spatially transformed inputs. We evaluated our method on several benchmark datasets, including a synthesised ImageNet test set. ITS outperforms the utilised baselines on all zero-shot test scenarios.
Canonicalisation, Geometric Deep Learning
1 Introduction
A primary objective in Deep learning is to develop models that exhibit robust generalisation during inference (commonly evaluated using test sets). To improve generalisation, models are trained with data sets of increasing variability and size. However, this often requires using larger models (Chefer et al., 2022). As novel datasets more closely mirror real-world scenarios, models undergo rigorous assessments in diverse, dynamic environments. This requires the models to learn intricate functions that generalise to all the variations presented in the data. Each new variation (dimension) scales the data space exponentially, known as the combinatorial explosion. Sampling in such spaces gets progressively harder (curse of dimensionality). For instance, consider an object from infinitely many perspectives. In a standard image classification task, all these variations have the same label assigned. Using 2D spatial transformations, these variations bring only redundant information. Hence, considering every conceivable variant is impractical, computationally infeasible, and even useless in terms of information gain. If a model can correctly predict an object from all perspectives, it is (pseudo-)invariant to these perspective changes.
Figure 1: Inspired by the human recognition process for perturbed inputs, we propose a novel inference framework coined Inverse Transformation Search (ITS). Similar to humans transforming the given input mentally or physically by tilting their heads or stepping forward, for instance. Instead of the usual likelihood estimation of a given sample, the pre-trained model can evaluate different input transformations before a final decision is made. The output of this process is the most familiar transformation of the query.
Data-driven Pseudo-Invariance
Over recent years, data augmentation (Hernández-García & König, 2018; Shorten & Khoshgoftaar, 2019) has become a routine instrument for increasing generalisation. The main drivers for this development are its simplicity and model agnosticism. Data augmentation enriches a dataset by applying transformations or modifications to existing data points. Augmented training data acts as an explicit Frobenius norm regularizer (LeJeune et al., 2019) promoting the extraction of consistent class features (Huang et al., 2021). Models either learn transformed replicates of the same filter (Olah et al., 2020) or fuzzy averages of transformed per-class features, regularising model complexity (Chen et al., 2019; Dao et al., 2019). This behaviour can be amplified using invariance-promoting losses (Shakerinava et al., 2022).
Goodfellow et al. (2009) showed that the quality of pseudo-invariance in the learned representation increases with the depth of the network. Minor deformations, like small shifts and rotations, are compensated in lower layers, while higher layers progressively form invariances. Augmented training can encourage this behaviour, as the model aims to align the kernel space with the largest eigenvector of the sample covariance (Balestriero et al., 2022b). However, strong forms of data augmentation, like cropping, can also hurt generalization as it can introduce feature biases by varying the dominance of features (Balestriero et al., 2022a).
Model-driven (Pseudo-)Invariance
Inductive biases for spatial symmetries are commonly developed for CNNs by exploiting their weight-shared filters (Schmidt & Stober, 2022). Instead of regularising the filter space using data, it can be constrained explicitly to symmetric weight matrices (so-called isotropic filters) (Hu et al., 2019). Invariance can also be achieved without such constraints by considering all transformed filter variations followed by a max-out (Gens & Domingos, 2014; Cohen & Welling, 2016; Bao & Song, 2020). This can be formulated more flexibly by a steerable basis using Bessel functions (Qiu et al., 2018; Delchevalerie et al., 2021) or 2D Hermite polynomials to allow for weight sharing across kernel orientations (Ecker et al., 2019; Ustyuzhaninov et al., 2020). However, these methods are limited to specific transformations, like centre rotations. Zhou et al. (2020) proposed to meta-learn the filter structure (binary mask) based on the equivariance required to solve a task.
Unfortunately, these inductive biases come with a significant engineering effort. The same applies to methods, which learn the geometry of the latent space (Arvanitidis et al., 2017) to find equivariance paths. With the introduction of Spatial Transformer Network (STN) (Jaderberg et al., 2015), another promising avenue began. These networks remove the spatial perturbations, i.e., canonicalise the transformed input signals by learning a highly non-linear mapping from pixel to transformation space. A more effective alternative is the mapping to canonical coordinate spaces (Esteves et al., 2018; Tai et al., 2019). To learn transformation compositions, STNs can be stacked (Lin & Lucey, 2017; Detlefsen et al., 2018). These transformations can also act on each pixel individually using pixel-wise motion fields (Shu et al., 2018). However, all of these variants require augmented training data. We propose a novel inference strategy that canonicalises inputs without augmented training.
Inference-driven Pseudo-Invariance
We take inspiration from human behaviour and propose a third way to integrate pseudo-invariances. The human brain learns to discard redundancies imposed by symmetries for a more efficient perception (Gómez et al., 2008; Leibo et al., 2011). To recognise a perceived object, our brain aligns it with our learned mental model of that object. This process is called canonicalisation (Graf, 2006) and it is either performed mentally or physically (or a mixture of both). We can leverage our physical abilities by squeezing our eyes, tilting our heads and slowly moving closer when triggered by an unclear perception. In this way, our trained mental model can properly recognise a variety of perturbed inputs by gradually aligning them to their known canonical form. This is a well-established theory supported by various experiments (Palmer et al., 1981; Harris et al., 2001; Gómez et al., 2008; Konkle & Oliva, 2011).
We can transfer this hypothesis by emulating the alignment search algorithmically as illustrated in Figure 1. We propose an Inverse Transformation Search (ITS) for trained classifiers during inference, which renders them zero-shot pseudo-invariant by only requiring access to their logit scores. We disentangle the process of class-specific feature learning (achieved during training) and canonical signal distillation (achieved during inference). The latter removes perturbations from the input and therefore eliminates variances from the query to ease the subsequent prediction.
Scope and Outline
In this work, we particularly focus on 2D image classifiers and the special linear group, comprising combinations of rotation, scaling, and shearing operations. However, our proposed method is more generally applicable. In Section 2, we start with an introduction to the language of symmetries, i.e., group theory. Upon these fundamentals, we build a group-induced evaluation score in Section 3 used in our search algorithm presented in Section 4. We evaluate our methods in Section 5 and conclude with future directions in Section 6.
2 Preliminaries
Understanding how to undo transformations requires a basic comprehension of symmetries and their mathematical formulations as groups. A symmetry of an object is a transformation that preserves the object’s structure. For instance, rotating an equilateral triangle around its centre by $0^{\circ}$, $120^{\circ}$ or $240^{\circ}$ leaves the triangle unaffected, as illustrated in Figure 2. These rotations form a group of finite size, i.e., a finite group.
Definition 2.1 (Group).
Groups are sets of objects equipped with a binary operation. For every group $G$ the following axioms must hold
$\begin{array}[]{ll}\text{closure:}&g,h\in G\rightarrow gh\in G,\\ \text{associativity:}&g(hk)=(gh)k,\\ \text{identity:}&eg=ge=g,\\ \text{invertibility:}&gg^{-1}=g^{-1}g=e,\end{array}$
with $g,h,k,e\in G$ and $e$ being the identity element.
This includes a composition of multiple group elements, like $gh$, which is the short form for $g\circ h$, where $\circ$ is some binary operator. Some groups possess a fifth property, that is, commutativity. We call them Abelian groups, e.g., centre-rotations. More details and proofs can be found in (Nash, 2016).
Let us extend the picture from discrete to continuous structures. For a circle, centre rotations of any angle send the object back to itself. Formally, these groups are called Matrix Lie groups (Hall, 2015), which form a smooth differentiable manifold. For the basic comprehension of this work, it is only important to understand that group elements can be represented by parameterised matrices. In this work, we focus on centre rotation, isotropic scaling, and shearing transformations (and their compositions). All of them are represented by $2\times 2$ matrices $\mathbf{T}$. This particular group is called the Special Linear Group $\text{SL}_{2}(\mathbb{R})$ (Hall, 2015).
When dealing with continuous structures, it is helpful to consider finite approximations instead. Given a group $G$ we can have a subgroup $G^{\prime}<G$, which still satisfies Def. 2.1 but with a smaller cardinality $|G^{\prime}|$. For instance, a subgroup of all centre-rotations is the finite group with multiples of $10^{\circ}$ rotations. Such groups can be generated by a single group element.
Definition 2.2 (Generator).
Given $g\in G$, $g$ can be used as a generator to form a finite (cyclic) subgroup $\{\langle g\rangle^{k}\mid k\in\mathbb{Z}\}$. $\langle g\rangle^{k}$ denotes $g$ to be a generator acting $k$-times.
A group element acts on an object $\mathbf{x}$ by transforming it. If the entire group $G$ acts on an object, the resulting output set is called the orbit of the object under $G$.
Definition 2.3 (Orbit).
Let $G$ be a group and $\mathbf{x}$ an object. Then, the orbit of $\mathbf{x}$ under $G$ is
$\text{orb}_{G}(\mathbf{x})=\{g\mathbf{x}\mid g\in G\}.$
All $g$ that leave $\mathbf{x}$ unchanged are symmetries of that object.
Definition 2.4 (Stabiliser).
The stabiliser of some $\mathbf{x}$ under $G$ is a subgroup that leaves $\mathbf{x}$ unchanged,
$\text{stab}_{G}(\mathbf{x})=\{g\in G\mid g\mathbf{x}=\mathbf{x}\}.$
We use stabilisers to formalise spatial symmetries, e.g., transformations that preserve the pixel positions of an image.
We can also formalise this from a function perspective. A function $f$ (e.g., modelled by a neural network), which captures a symmetry defined by group $G$, is either equivariant, invariant, or covariant (Marcos et al., 2017).
Definition 2.5 (Variances).
Let $f(\mathbf{x})$ be a function of an input $\mathbf{x}$. Then, $f$ is
$\begin{array}[]{llll}\text{equivariant}&\text{iff}&f(g\mathbf{x})=gf(\mathbf{x}),&\forall g\in G,\\ \text{invariant}&\text{iff}&f(g\mathbf{x})=f(\mathbf{x}),&\forall g\in G,\\ \text{covariant}&\text{iff}&f(g\mathbf{x})=g^{\prime}f(\mathbf{x}),&\forall g\in G,g^{\prime}\in G^{\prime},\end{array}$
with $G^{\prime}$ being isomorphic to $G$.
Note that covariance is the most general one, where both transformations differ in their domain and co-domain, respectively. Invariance and equivariance are special cases of covariance. If Def. 2.5 does not hold but is approximately equal, we call the variance a pseudo-variance. For more elaborated details about variances in deep learning, see (Bronstein et al., 2021).
3 Localizing Canonical Forms
In this section, we introduce a novel method for finding canonical forms. We use $f_{\mathbf{\theta}}:\Omega\to\mathbb{R}^{Y}$ to denote a parameterised function mapping from pixel space to class-logits. We call this function a softmax-classifier if it parameterises a categorical distribution by its logits, such that
$p_{\mathbf{\theta}}(y\mid\mathbf{x})=\operatorname{softmax}\left(f_{\mathbf{\theta}}(\mathbf{x})\right):=\frac{\exp\left(f_{\theta}(y\mid\mathbf{x})\right)}{\sum_{y^{\prime}}\exp\left(f_{\theta}(y^{\prime}\mid\mathbf{x})\right)}.$ (1)
During training, we obtain the posterior
[…]
$G$-transformed Test Sets
Commonly, models are tested on holdout sets with the same spatial variation. Instead, we aim to train a classifier on canonical examples (vanilla training data) and subsequently evaluate it on a $G$-transformed test set. This transformation introduces a distribution shift between $\mathcal{D}_{\text{train}}$ and $G\mathcal{D}_{\text{test}}$. Therefore, requiring $p(y\mid\mathbf{x},\mathcal{D}_{\text{train}})$ to perform a zero-shot task on the out-of-distribution test set. We argue, that this more accurately reflects real-world scenarios. Under slight abuse of notations, we define $G\mathcal{D}_{\text{test}}\coloneqq\{\langle\mathbf{x}^{\prime},y\rangle\mid\mathbf{x}^{\prime}=g\mathbf{x},\langle\mathbf{x},y\rangle\sim\mathcal{D}_{\text{test}},g\sim G\}$. Note that, $G\mathcal{D}_{\text{test}}$ and the vanilla $\mathcal{D}_{\text{test}}$ intersect as the stabilisers of all samples are contained. For all our experiments, we use $\mathcal{D}_{\text{train}}$ for training and $G\mathcal{D}_{\text{test}}$ during inference.
Figure 2: (left) Centre-Rotating an equilateral triangle and (right) an exemplary confidence surface.
Max-Confidence Estimation
We derive an energy-based evaluation from the concept of invariances. We assume that the function $f_{\mathbf{\theta}}$ is sensitive to $G$, which is true for all neural networks when $G$ is getting more complex. That is, for any non-trivial element $g$ in $G$ we have $f_{\mathbf{\theta}}(\mathbf{x})\neq f_{\mathbf{\theta}}(g\mathbf{x})$. However, based on Def. 2.1, there exists a $g^{-1}\in G$ such that $f_{\mathbf{\theta}}(\mathbf{x})=f_{\mathbf{\theta}}(g^{-1}g\mathbf{x})$. Using a confidence measure $\rho:f_{\mathbf{\theta}}(\cdot)\to\mathbb{R}$ (Kaba et al., 2023), we can determine $g^{-1}$ by
$g^{-1}\leftarrow\text{arg}\max_{h\in G}\rho\left(f_{\mathbf{\theta}}(hg\mathbf{x})\right).$ (3)
This would allow us to undo perturbations induced by real-world measurements. An illustration of a confidence surface is depicted in Figure 2 (right). As $\rho$ is the limiting factor of such an approach, we study three different options in the following.
Reducing Prerequisites
Any such confidence measure relies on a semantically structured latent space (Bengio et al., 2012). Although a multitude of methods exist to access prediction confidence. Many of these either pose a prior on the model to learn confidence during training (Moon et al., 2020) or require access to the training data to approximate the data manifold (Snell et al., 2017; Kotelevskii et al., 2022). To maintain a minimal level of obstacles, we seek a post-hoc method without any knowledge about the training set.
Energy-induced Confidence
The probability masses of softmax-classifiers are often spoiled by overconfident posteriors even for out-of-distribution samples (Lee et al., 2018; Geifman et al., 2018). Therefore, the model’s energy surface is often used as a straightforward approximation to the confidence surface (Liu et al., 2020). Duvenaud et al. (2020) showed that Energy-based models (LeCun et al., 2006) are implicitly contained in both discriminative and generative models. This renders Energy-based measures widely applicable. Following Duvenaud et al. (2020) we determine the total energy of a model $f_{\mathbf{\theta}}$ by
$E_{\mathbf{\theta}}(\mathbf{x})=-f_{\mathbf{\theta}}(\mathbf{x})=-\log\sum_{y}\exp f_{\mathbf{\theta}}(y\mid\mathbf{x}),$ (4)
with $p_{\mathbf{\theta}}(\mathbf{x})\propto\sum_{y^{\prime}}\exp\left(f_{\theta}(y^{\prime}\mid\mathbf{x})\right)$. We denote class-wise energies by $E_{\mathbf{\theta}}(y\mid\mathbf{x})=-f_{\mathbf{\theta}}(y\mid\mathbf{x})$. Using this relation, we can use any trained classifier and interpret its logit scores as energies. Liu et al. (2020) showed that the negative energy is a straightforward confidence estimate.
Bayesian-induced Confidence
Commonly, logits in Equation 1 are parameterized by a single $\mathbf{\theta}$ obtained by Maximum Likelihood Estimation on $\mathcal{D}_{\text{train}}$. Instead of such point estimates, Bayesian neural networks (Osband, 2016) approximate a parameter distribution, hence enabling richer statistical insights, such as confidence estimation. Unfortunately, exact posterior estimation is known to be intractable due to the integration over the parameter space. Unlike common approximation schemes (Osawa et al., 2019) including ensemble methods (Lakshminarayanan et al., 2017), we only have access to a trained classifier to ease the applicability of ITS.
To the best of our knowledge, the only method to fulfil this requirement is Monte-Carlo Dropout (Gal & Ghahramani, 2016). Each parameter is endowed with a random process regulating its influence in the forward pass during inference. Let $\mathbf{\Lambda}_{l}:=\operatorname{diag}(\boldsymbol{\lambda}_{l})\sim\text{Bernoulli}(\lambda\in[0,1])$ be a diagonal matrix of random variables for some layer $l$. For each inference pass, we sample $\mathbf{\Lambda}_{l}$ and apply it to the weights as multiplicative noise, $\mathbf{W}_{l}\leftarrow\mathbf{\Lambda}_{l}\mathbf{W}_{l}$. We can steer the chance of dropout by $1-\lambda$. This allows us to define an unbiased estimator for the expectation,
[…]
The model’s confidence corresponds to the negative total uncertainty, which we can estimate using the entropy,
$\mathbb{H}\left[\mathbb{E}_{p_{\mathbf{\theta}}(y\mid\mathbf{x})}\left[y\right]\right]\approx-\sum_{y}\mathbb{E}_{p_{\mathbf{\theta}}(y\mid\mathbf{x})}\left[y\right]\log_{2}\mathbb{E}_{p_{\mathbf{\theta}}(y\mid\mathbf{x})}\left[y\right].$ (6)
Figure 3: Probability of recognising a transformed sample to be the canonical form. Each evaluation point represents the mean over the FashionMNIST test set.
Confidence Anomalies
The confidence measure should concentrate all its mass around the canonical form. However, we found that both baselines tend to increase their mass towards the domain borders. This is best observed in the rotation experiment in Figure 3 (last row). Here, the spatial alignment bias of FashionMNIST (Xiao et al., 2017) leads to increased confidence for $90^{\circ}$ rotated images (high activation of horizontal edge detectors). Similar confidence bumps appear when scaling up or down a sample until pixelation or disappearance. We propose a more stable measure, which does not suffer from these confidence anomalies.
Group-induced Confidence
Without explicit training or inductive biases, neural nets corrupt the geometry of group orbits during encoding (Arvanitidis et al., 2017). When traversing the orbit’s trajectory, the energy will fluctuate (Li et al., 2018). As both confidence scores utilise the energy, the confidence will fluctuate likewise. We remove noise from the energy by estimating its expected value $\bar{E}_{\mathbf{\theta}}(\cdot)$ using Monte-Carlo Dropout. This renders the measure stochastic. We use a generator $\hat{g}$ to generate the finite group $G$. We can evolve $\mathbf{x}$ along $\text{orb}_{G}(\mathbf{x})$ by incrementing or decrementing $n$ in $\hat{g}^{n}\mathbf{x}$.
Instead of evaluating each point independently, we consider the behaviour in the local neighbourhood as well. We use the curvature of the energy along the orbit. Local maxima have high negative curvature, which we approximate by the second derivative. This leads to the confidence measure,
$\rho_{G}(\mathbf{x},g)=-\frac{\partial^{2}}{\partial g^{2}}\big{(}\bar{E}_{\mathbf{\theta}}\star K\big{)}\big{(}g\mathbf{x}\big{)},$ (7)
where $K$ is a Gaussian smoothing kernel to reduce local over-confidence. To ensure that the confidence decreases at domain borders, we use “nearest” padding. In practice, we can approximate Equation 7 as follows.
Theorem 3.1 (Group Confidence).
Let $\phi_{\bar{E}_{\mathbf{\theta}},K}(\mathbf{x},g):=\big{(}\bar{E}_{\mathbf{\theta}}\star K\big{)}\big{(}g\mathbf{x}\big{)}$. The group-induced confidence defined in Equation 7 can be approximated by
$\rho_{G}(\mathbf{x},g)=-\lim_{|G|\to\infty}\frac{1}{\|\hat{g}-\hat{g}^{2}\|^{2}}$ (8)
$\bigg{(}\phi_{\bar{E}_{\mathbf{\theta}},K}(\mathbf{x},\hat{g}\circ g)+\phi_{\bar{E}_{\mathbf{\theta}},K}(\mathbf{x},\hat{g}^{-1}\circ g)-2\phi_{\bar{E}_{\mathbf{\theta}},K}(\mathbf{x},g)\bigg{)}.$
Find the proofs in Section D.1 and Section D.2.
4 Inverse Transformation Search
Figure 4: Our proposed Inverse Transformation Search (ITS) on a perturbed image $\hat{\mathbf{x}}$. In each level of the search tree, the orbit of one $G$ is evaluated. At each level all samples are evaluated using Equation 7 and a pre-trained model $f_{\theta}$.
Figure 5: An example iteration of our algorithm with a generated sub-group of size $9$ and $2$ hypotheses. The first layer rotates the input, the second layer scales it, and the last layer shears it.
An overview and an example of our proposed search procedure are illustrated in Figure 4. In the following, we spell out the most important aspects of it with more details in Appendix C.
Transforming Images
An image $\mathbf{x}$ is a collection of pixel vectors containing spatial and colour information. Spatial transformations are modifications of the location a pixel vector is pointing to. To resample the transformed image, colour values are interpolated between the modified pixel positions. This procedure can be efficiently executed in parallel on the GPU (Jaderberg et al., 2015).
Forming Search Levels
We assume that a vanilla data point $\mathbf{x}$ is transformed by some groups acting on it. Let $\hat{\mathbf{x}}:=(g_{0}\circ g_{1}\circ\ldots\circ g_{m})\mathbf{x}$ where $g_{0},g_{1},\ldots,g_{m}$ might be elements of different groups. We follow a heuristic (greedy) approach by decomposing the transformation into different groups acting sequentially on the input. This reduces the complexity of the problem significantly. We propose an iterative search procedure, which aims to eliminate one $g_{k}$ at every step using its inverse $g_{k}^{-1}$. This assumes either that the order of transformation is known, which is unlikely in real-world settings or that we deal with Abelian groups. We show that the restrictions imposed by the latter can be attenuated, as only a sub-sequence needs to be Abelian.
Theorem 4.1 (Group Elimination).
Let $s=g_{0}\circ g_{1}\circ\ldots\circ g_{m}$ be a finite sequence of group elements with $g_{0}\in G_{0},g_{1}\in G_{1},\ldots,g_{m}\in G_{m}$ and $m\in\mathbb{N}$. Further, let $g_{k}^{-1}$ be the inverse of $g_{k}$ in the sequence for some $k\in[0,m]$ and $s_{\setminus g_{k}}$ being $s$ without $g_{k}$. We have, $g_{k}^{-1}s=s_{\setminus g_{k}}$ if all elements in the sub-sequence $s_{0:k-1}$ and $g_{k}^{-1}$ are elements of an Abelian group.
The proof can be found in Section D.3. To cope with Lie groups, we generate sub-groups of size $|G|$ using generators. In that light, each layer comprises $|G|$ elements (as a finite set of equidistant samples along the orbit). This allows us to express transformations by an exponential $\hat{g}^{n}$ with $n\in\mathbb{N}$. This renders the evaluation in each layer tractable, as only a finite set of points needs to be evaluated. [^1] All non-cyclic groups, like translations, would at some point shift the image signal outside the visible pixel grid. In practice, we consider only a reasonably bounded set of transformations.
Figure 6: Evolution of $H_{0}$ and $H_{1}$ in a two-layer search tree (rotation in the top row and shearing in the bottom one) on Fashion-MNIST. Red areas indicate an average superiority of $H_{0}$, while blue areas indicate a superiority of $H_{1}$.
Hypothesis Testing
We define a hypothesis $H_{i}$ as a trajectory down the search tree. Humans use selective hypothesis testing (Sanbonmatsu et al., 1998) when the perceived visual stimulus has high uncertainty. This includes evaluating multiple hypotheses until a final decision is made. We simulate this behaviour by traversing the search tree $k$ times in parallel. This is done by updating a collection of candidate hypotheses $\{H_{0},H_{1},\ldots,H_{k-1}\}$. After the first expansion and evaluation step, we select the $k$ best candidates ordered by confidence. Each further expansion and evaluation step is done for all $H$. The $k$ best are selected from the cumulative candidate set, bounding each level of the tree to $k$ children. Find an example of such a $k$-ary tree in Figure 5.
Rather than assuming independent trajectories, we pose a linking condition between them to avoid redundant hypotheses. For instance, a rotated image of a $5$ might share many features with a $9$. If the confidence is high, a significant amount of runs would refine this hypothesis until the end, resulting in a one-sided perspective. We improve the hypothesis testing by allowing only one hypothesis per label to be further refined.
Change of Mind
In theory, such a dynamic search process over multiple hypotheses would allow the model to ”change its mind”. That is, rejecting $H_{0}$ (the initially dominating hypothesis) and instead accepting some other $H$ for the final incumbent (output of the inference process). Figure 5 already provides an example for such phenomena, which we investigate more thoroughly using Figure 6. We constructed a regular grid of rotated and sheared Fashion-MNIST test sets. Each point in Figure 6 is the average performance of a CNN with ITS (see Section B.2 for more details). Colour changes from layers 1 to 2 indicate that the initial hypothesis is on average superseded by the other. To maximise the performance outcome (we focus on classification accuracy), we need to substitute the leading hypothesis $H_{0}$ with its alternative $H_{1}$ in some regions. Without a ground-truth signal (during test time), we need a measure to determine when to swap hypotheses. In our experiments, we used class support, which is the number of occurrences of the predicted class over the entire tree. However, this indicator has only minor advantages over the confidence score (see Figure 6 e,f). We leave this promising avenue for future research.
Complexity and Runtime
We have a fixed-width $k$-ary tree, where the number of elements in the first level is $n:=|G|$ and in all consecutive levels $kn$. As all evaluations are performed in parallel, $n$ does not affect the runtime. For our implementation, the Monte Carlo estimates are performed sequentially scaling the complexity by $M$. This gives $\mathcal{O}(Mk^{d})$, where $d$ denotes the depth of the tree. We quantified the runtime cost in Section B.1. Our inference strategy is particularly interesting for applications where it is worth trading runtime against robustness (Ghamizi et al., 2022).
5 Experiments
In each classifier, there is potential for pseudo-invariance, which we just have to activate. In that light, all investigations are zero-shot performance tests with out-of-distribution data (see Section 3). We split the vanilla datasets into disjunct training, validation, and test sets. We always employ the vanilla training set to fit the model and validate it on the vanilla validation set. All performance measures are subsequently evaluated on the perturbed test set $G\mathcal{D}_{\text{test}}$, where $G$ is specified in each experiment.
If not further specified, we used zero-padding to define areas outside the pixel space $\Omega$, bilinear interpolation, and a group cardinality of $n=17$. We used two CNN-backbones simple as the LeNet architecture (Lecun et al., 1998) (see Section A.1). All models are trained with the AdamW optimizer (Loshchilov & Hutter, 2017) using default parameters. We minimised the negative log-likelihood using ground-truth image labels. We used a learning rate of $35e-$ 3, $3$ epochs for MNIST, $5$ epochs for Fashion-MNIST and mini-batches of size $128$. In all MNIST experiments, we removed the class $9$ samples, as this introduces a class-superposition for $180^{\circ}$ rotations. You can find more details in our publicly available source code. [^2] Added after blind reviews.
5.1 Zero-Shot rotated Image Classification
Figure 7: Test accuracy on rotated MNIST with fixed angles.
We trained a CNN, a GCNN (Cohen & Welling, 2016) and a RotDCF (Cheng et al., 2018) on the vanilla (canonical) MNIST. Both baselines use Filter Decomposition to gain $p4$-equivariance and $p8$-equivariance, respectively. We evaluated their zero-shot performances on the rotated test sets with results shown in Figure 7. Such a model would be robust to centre-rotations. Just by exploiting the confidence surface of the vanilla (shallow) baseline CNN our method boosts its robustness significantly. One can also see, that the rotation-equivariant GCNN outperforms our methods in the majority of test angles. However, such models require hard-coded inductive biases tailored to some specific groups [^3] Developing inductive biases for all possible transformations is impossible due to the vastness of possible transformations.. Our method does not pose any such requirements in the hypothesis space, also allowing for smaller model complexity. We enable pseudo-equivariance to more transformations, which is shown in the subsequent experiments.
5.2 Zero-Shot Affine Canonicalisation
Figure 8: Three test set samples of MNIST (top), Fashion-MNIST (middle), GTSRB (Stallkamp et al., 2012) (bottom), which are transformed by a composed spatial transformation of rotation $[-\pi,\pi]$, scaling $[-\frac{1}{4},\frac{1}{4}]$ and shearing $[-\frac{1}{4},\frac{1}{4}]$. All models use the same CNN backbone, are trained on $\mathcal{D}_{\text{train}}$ and are tested on $G\mathcal{D}_{\text{test}}$. ITS is used with $3$ and $5$ levels and compared to a STN and a DTN with a single and $3$ transformation layers, respectively.
Table 1: Classification accuracies ($\uparrow$) and Frobenius distances ($\downarrow$) between the ground truth and the predicted transformation matrix. Mean and standard deviations are estimated over $25$ test set runs with random rotation, scaling, and shearing parameters.
Dataset CNN STN STN3 DTN ETN ITS3 ITS5
MNIST 50.15$\pm$9.29 49.78$\pm$9.13 13.85$\pm$4.05 48.61$\pm$8.99 56.59$\pm$9.87 88.89$\pm$13.86 89.81$\pm$13.95
F-MNIST 25.91$\pm$5.93 24.51$\pm$5.81 12.42$\pm$3.73 24.11$\pm$5.71 26.10$\pm$5.98 38.47$\pm$7.77 37.58$\pm$7.56
GTSRB 32.01$\pm$6.59 31.46$\pm$6.55 2.78$\pm$1.69 28.21$\pm$6.23 40.02$\pm$7.99 66.44$\pm$11.48 67.09$\pm$11.53
MNIST 23.03$\pm$2.48 25.43$\pm$2.72 23.61$\pm$2.54 31.02$\pm$3.31 26.10$\pm$2.83 18.80$\pm$2.22 18.98$\pm$2.32
F-MNIST 23.03$\pm$2.80 25.96$\pm$3.14 23.61$\pm$2.87 35.16$\pm$4.22 34.02$\pm$4.19 20.25$\pm$2.59 20.30$\pm$2.56
GTSRB 23.04$\pm$2.85 18.15$\pm$2.23 23.62$\pm$2.92 34.61$\pm$4.23 21.08$\pm$2.61 17.87$\pm$2.26 17.93$\pm$2.27
Table 2: Average acc@1 and acc@5 over the SI-Rotation test set using a ViT-B16 pre-trained on ImageNet. The vanilla performance is compared to a rotation and scale augmented fine-tuning, training on foreground objects with and without annotated segmentation masks (AS) and our ITS. We used two acc@5 estimates for ITS: over $H_{0}$’s categorical distribution and over the top-1 predictions of all $H$.
SI-Rotation Test Set Samples ViT acc@1 acc@5
[Uncaptioned image] vanilla 38.5 58.5
Rotation + Scale Augm. 41.0 (+2.5) 63.2 (+4.7)
(Chefer et al., 2022) 44.8 (+6.3) 65.2 (+6.7)
(Chefer et al., 2022)+AS 46.2 (+7.7) 67.0 (+8.5)
ITS (ours) 49.0 (+10.5) 68.7 (+10.2)
70.9 (+12.4)
Similar to STNs, we can undo affine transformations (i.e., canonicalize). We compare ITS to a STN (Jaderberg et al., 2015) with one and three transformation layers using inverse compositional layers (Lin & Lucey, 2017), a Diffeomorphic Transformer Network (DTN) (Detlefsen et al., 2018), and a Equivariant Transformer Network (ETN) (Tai et al., 2019). To study the impact of depth, we contrasted a $3$- and a $5$-layer ITS. We used a layer stack of rotation + scaling + shearing + (rotation + scaling). We refer to these as ITS 3 and ITS 5, respectively. Table 1 shows the quantitative results of our experiments. The associated qualitative results as reconstructions (after applying the predicted inverse transformation) can be found in Figure 8.
We found all baselines incapable of performing affine canonicalisations outside the training data distribution. This makes sense, as their localisation networks (regressing the transformation parameters) learn pixel-to-transformation mappings from training data. They are unable to generalise - particularly to rotations, which are mostly present during testing. The STN and the DTN estimate the composed transformation matrix in one forward pass. They need to extrapolate from nearly no rotations during training to $360^{\circ}$ rotations during testing. As the rotation manifold is isomorphic the unit circle ($\cos(\theta),\sin(\theta)$), this non-linearity needs to be learned implicitly. The ETN on the other hand uses canonical coordinates, which linearise the relationship between rotations.
While ITS does not always lead to correct results, it works best among the contestants. In Table 1 we also see that for most cases the Frobenius distance is lowest, too. Interestingly, the distance to the ground truth matrix is lowest for ITS 5 while ITS 3 achieves higher accuracies. High Frobenius distance and high accuracy would indicate that the classifier got an incorrectly rectified sample but is still able to predict the correct class.
5.3 Scaling up to ImageNet
Si-Score (short SI) (Djolonga et al., 2021) is a synthetic vision dataset for robustness testing, comprising semantically masked ImageNet (Russakovsky et al., 2015) objects pasted on various backgrounds. The foreground objects are either arbitrarily (horizontally and vertically) translated (SI-Location), arbitrarily rotated (SI-Rotation), or arbitrarily scaled (SI-Size). We evaluate ITS using a Vision Transformer (ViT)-B16 (Dosovitskiy et al., 2021). The Vision Transformer is pre-trained and validated on ImageNet and tested on Si-Score. We trained the model with and without data augmentation using random rotations $[-\pi,\pi]$ and scales $[-\frac{1}{4},\frac{1}{4}]$.
We found significant improvements on SI-Rotation as denoted in Table 2. To some extent this was expected as our confidence measure is highly sensitive in detecting canonical rotations (see Figure 3). Surprisingly, however, our method even outperformed augmented fine-tuning and explicit training on the foreground object (Chefer et al., 2022). On SI-Size ITS maintains the performance of the baseline. We hypothesise that location and size are properties that have high variance in the vanilla ImageNet already. This puts the test performance on a high level without requiring any additional mechanisms. Equipping the backbone with the ability to translate the input query with two degrees of freedom causes it to focus on the wrong object. This behaviour is similar to attention but with a fixed centre bias of the backbone and a spatial translation of the signal instead. However, ITS does not get any feedback (gradient) if the object of focus is correct. This leads to slightly degenerated performance on SI-Location (see Section B.3).
6 Conclusion
Inverse Transformation Search (ITS) equips any vanilla image classifier with pseudo-invariance to various spatial transformations during inference. This increases its robustness to spatially perturbed inputs without requiring extra training data. We activate this potential, which lies hidden in any classifier, without requiring any inductive biases or augmented training. This is achieved by an incremental search for the inverse perturbation. At the canonical form, the features of the transformed input match the features known from the training data. Here the model is most confident to classify this sample properly. To mitigate the issue of spatial collapses during repetitive transformations, we propose a group-induced confidence measure. We successfully showed the superiority of our methods for zero-shot image classification tasks.
Future Work
One issue is the rigidness of existing image datasets. Heavily perturbed images can have multiple proper hypotheses. See the example in Figure 6, where both labels ($5$ or $9$) seem reasonable. As ITS uses the representation space of the trained classifier, its performance depends on the quality of these representations. The model might detect features in transformed signals, which lead to false hypotheses. Even worse, these hypotheses might be reinforced down the search tree. Easing these instabilities should be the main objective of future investigations.
7 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Implementation Details
For educational purposes, we used $M=1$ (i.e., a deterministic search), which ensures consistently improving scores along the paths in Figures 5 and 6. We used a group cardinality of $n=17$ for all our experiments due to algorithmic reasons. We divide the parameter space $[-\theta,\theta]$ by $n$ into an equally distant parameter set. The identity $\theta=0$ lies in the middle of this domain and is therefore only included for odd $n$. For instance, $n=7$ gives $3$ clockwise rotations, $3$ counter-clockwise rotations and the identity (no rotation).
A.1 Model Architectures
Small-Sized CNN
This is used for experiments on MNIST and Fashion-MNIST. This model comprises three 2D-convolutional layers with $5\times 5$ kernels and $8$, $16$, and $16$ channels. We add GeLU non-linearities (Hendrycks & Gimpel, 2023) after each convolution layer. 2D Max-Pooling with $2\times 2$ kernels is added after the second and the third convolutional layer. The final stack of feature maps is flattened into a vector and further compressed to the number of classes by two linear layers. After the first linear layer, we also add a GeLU non-linearity and a dropout-layer with $50\%$ of dropout. From here we read out the logit scores.
Large-Sized CNN
This is used for experiments on GTSRB. This model comprises three 2D-convolutional layers with kernel sizes of $5$, $3$ and $1$ and $100$, $150$, and $250$ channels. We add GeLU non-linearities (Hendrycks & Gimpel, 2023), 2D Max-Pooling with $2\times 2$ kernels, Batch Normalisations (Ioffe & Szegedy, 2015) and Dropout after each convolution layer. The final stack of feature maps is flattened into a vector and further compressed to the number of classes by two linear layers. After the first linear layer, we also add a GeLU non-linearity and a dropout-layer with $50\%$ of dropout. From here we read out the logit scores.
Appendix B Further Experiments
B.1 Runtime
Figure 9: Walltime of ITS (shades of blue) with a CNN backbone (gray) on a transformed MNIST test set (standard deviation shown as error bars).
All experiments are performed on an Nvidia A40 GPU (48GB) node with 1 TB RAM, 2x 24core AMD EPYC 74F3 CPU @ 3.20GHz, and a local SSD (NVMe). The software specifications of our implementations can be found in our open-sourced code. Figure 9 shows how the runtime behaves under different hyper-parameter settings compared to the pure backbone. The most significant factor is the depth of the search tree. In Section 5.2 we found that the performance does not notably increase with more hypotheses. The number of Monte-Carlo dropout steps, $M$, stabilises the predictions by reducing temporal confidence anomalies. Although, we used $M=50$ for our experiments, the impact on the performance is minor. Another factor is the number of hypotheses tested. These hypotheses are processed in parallel but due to the unique class condition (see Section 4), the runtime scales with $k$. Empirically, we found that $k=2$ produces the best results, which we used in all of our experiments. Furthermore, we want to emphasise again that our inference strategy is targeted for applications where it is worth trading runtime against robustness, as in (Ghamizi et al., 2022).
B.2 Change of Mind
To compute the results in Figure 6, we used the following measures: We estimated the transformation distance between the predicted matrix, $\mathbf{T}_{H_{0}}$ or $\mathbf{T}_{H_{1}}$, and the ground-truth matrix, $\mathbf{T}_{\mathrm{GT}}$. Let $d(\mathbf{T}_{H_{i}},\mathbf{T}_{\mathrm{GT}}):=\|\mathbf{T}_{H_{i}}-\mathbf{T}_{\mathrm{GT}}\|_{F}$ for $i\in\{0,1\}$. To ease comparability, we used a cross-normalisation to relate the outcomes. Let $\mathcal{T}_{G\mathcal{D}_{\text{test}}}[H_{i}]$ be the set of predicted transformation matrices of $H_{i}$ for all samples in $G\mathcal{D}_{\text{test}}$. Then, the minimum and maximum distance can be obtained by
$d_{\text{min}}$ $:=\min_{\mathbf{T}\in\mathcal{T}_{G\mathcal{D}_{\text{test}}}[H_{0}]\cup\mathcal{T}_{G\mathcal{D}_{\text{test}}}[H_{1}]}d(\mathbf{T},\mathbf{T}_{\mathrm{GT}}),$ (9)
$d_{\text{max}}$ $:=\max_{\mathbf{T}\in\mathcal{T}_{G\mathcal{D}_{\text{test}}}[H_{0}]\cup\mathcal{T}_{G\mathcal{D}_{\text{test}}}[H_{1}]}d(\mathbf{T},\mathbf{T}_{\mathrm{GT}}).$ (10)
We can min-max cross-normalise each distance by
$\bar{d}(\mathbf{T}_{H_{i}},\mathbf{T}_{\mathrm{GT}})=\frac{d(\mathbf{T}_{H_{i}},\mathbf{T}_{\mathrm{GT}})-d_{\text{min}}}{d_{\text{max}}-d_{\text{min}}}\in[0,1].$ (11)
We are interested in the difference between both predictions,
$\bar{d}(\mathbf{T}_{H_{0}},\mathbf{T}_{\mathrm{GT}})-\bar{d}(\mathbf{T}_{H_{1}},\mathbf{T}_{\mathrm{GT}}).$ (12)
This is the measure used to produce each point in Figure 6 (a,c). The construction of the classification accuracy plots in Figure 6 (b,d) is done similarly. We counted the correct top-1 class-predictions of $H_{0}$ and $H_{1}$ over $G\mathcal{D}_{\text{test}}$. Again, cross-normalised the results and computed the difference.
The confidence score differences and the class support differences are only plotted for the last layer. This is done as the re-ranking of the hypotheses is only required at the end of the search procedure for the final read-out of the incumbent. Both measures are computed similarly to the previous two. The confidence scores and the class support counts are compared by relational operators and the result is stored as binary matrices. We sum their entries, normalise and compute the difference as before.
The idea of the class support is to identify the hypothesis where the associated class label appears most often during the search. Therefore, most transformed variations fall into one class cluster within the representation space. We argue that this might be evidence for the correctness of the prediction. Also, an equivariant model would maximise this measure and assign all samples to one class.
B.3 Continuation of Section 5.3
Test Set Samples ViT acc@1 acc@5
[Uncaptioned image] vanilla 54.9 77.5
Rotation + Scale Augm. 45.3 (-9.6) 67.9 (-9.6)
(Chefer et al., 2022) 60.2 (+5.3) 80.6 (+3.1)
(Chefer et al., 2022)+AS 61.0 (+6.1) 81.4 (+3.9)
ITS 55.1 (+0.2) 77.7 (+0.2)
72.3 (-5.2)
[Uncaptioned image] vanilla 33.3 52.2
(Chefer et al., 2022) 38.6 (+5.3) 57.0 (+4.8)
(Chefer et al., 2022)+AS 38.6 (+5.3) 57.8 (+5.6)
ITS 29.7 (-3.6) 45.0 (-7.2)
43.6 (-8.6)
Table 3: Continuation of Table 2 with SI-Size and SI-Location.
Figure 10: SI-Location sample showing a pizza (ground-truth label) piece on a lakeside background. In the first layer, $H_{0}$ classified the image as a lakeside and $H_{1}$ as a pizza. In the second layer, $H_{0}$ got refined but $H_{1}$ changed its hypothesis to now being a drilling platform.
Table 3 shows the continuation of our experimental results presented in Table 2. On SI-Size and SI-Location ITS is not able to push the performance of the baseline ViT. We found no improvement on both test sets. One driving factor is the ability of ITS to focus on different parts of the image if translation is used in at least one of the layers. This ”issue” is reinforced when another layer can scale the signal. This creates a magnifying glass effect on objects, that the backbone networks predicts are the target objects. As ImageNet classes are diverse, features might appear in the background leading the network to focus here. In Figure 10 we provide an example of a focus failure.
Appendix C Theoretical Specifications
C.1 Transformation Decomposition
Group representations of $\operatorname{Aff}_{2}(\mathbb{R})$ are defined as $3\times 3$ matrices with real entries (Hall, 2015). These matrices can be decomposed into a centre-rotation, isotropic scaling, reflection, shearing and translation matrix of the same size. We use these basic (or atomic) transformation sub-groups to define the orbits of each search level.
Centre-rotations and isotropic scaling can be represented by a single parameter. Therefore, these sub-groups can be generated by a single generator. For the remaining sub-groups, two generators are required. Reflection is either done about the horizontal or vertical centre axis. Shearing and translation are further decomposed as follows: Either of these transformations is performed along and against the direction of some vector positions at the centre.
C.2 Convolution over Finite Orbits
With this decomposition, convolution over the orbit of any such sub-group can be defined. This is essential for our group-induced confidence measure in Equation 7.
Lemma C.1 (Convolution over Finite Orbits).
Convolving some function $f$ with the orbit, $\operatorname{orb}_{G}(\mathbf{x})$, can be expressed by
$\left(\operatorname{orb}_{G}(\mathbf{x})[n]\star f\right)(g)=\sum_{m}\operatorname{orb}_{G}(\mathbf{x})[m]f((\hat{g}^{n})^{-1}\hat{g}^{m}).$
C.3 Markov Interpretation
Each branching candidate $g^{-1}$ in the search process is a result of Equation 3. All of the aforementioned measures (see Equations 4, 6 and 7) are stochastic when simulating Bayesian inference (as we do using Monte-Carlo dropout). The evaluations are therefore based on a random variable, which allows us to formulate the process of traversing the search tree as a stochastic process.
More specifically, the search describes a Markov process. The Markov property of a stochastic process guarantees that the probability of a future state depends only on the current state independently of its past behaviour. This implies that the branching candidate in each level only depends on the previous candidate. This is true, as every roll-out depends only on the prior candidate in the first place. These one-step dynamics simplify the system. In Section D.4 we provide a sketch of the proof.
Appendix D Proofs
D.1 Proof: Lemma C.1
Proof.
Given some kernel $K$ over the orbit of some finite group, $G$. Let $\operatorname{orb}_{G}(\mathbf{x})$ be the orbit of $\mathbf{x}$ under the action of $G$. Convolving $\operatorname{orb}_{G}$ with $K$ is per definition given by
[…]
As $G$ is a finite group, we can simplify the integration by the Riemann summation over $G$,
$\left(\operatorname{orb}_{G}(\mathbf{x})[g]\star K\right)(g)=\sum_{h\in G}\operatorname{orb}_{G}(\mathbf{x})[h]K(g^{-1}h).$ (14)
This can be rewritten using generators, such that
$\left(\operatorname{orb}_{G}(\mathbf{x})[n]\star K\right)(g)=\sum_{m}\operatorname{orb}_{G}(\mathbf{x})[m]K((\hat{g}^{n})^{-1}\hat{g}^{m}).$ (15)
It is to be shown that $\operatorname{orb}_{G}(\mathbf{x})$ can be evaluated by $n\in\mathbb{N}$ instead of $g$. This needs to be justified for the single and dual generator case.
Single Generator
For cyclic groups, where one generator suffices, every $g\in G$ can be substituted by the generator raised to some power. The orbit map from Def. 2.3 can be rewritten as
$\operatorname{orb}_{G}(\mathbf{x})$ =$=\{g\mathbf{x}\mid g\in G\}$
=$=\{\hat{g}^{n}\mathbf{x}\mid n\in[0,1,\ldots,|G|-1]\}.$ (16)
where $\hat{g}^{n}$ denotes the generator raised to power $n$. This isomorphism between $g$ and $n$ also simplifies the evaluation
$\operatorname{orb}_{G}(\mathbf{x})[g]=\operatorname{orb}_{G}(\mathbf{x})[n].$ (17)
Dual Generator
The second case assumes, that $G$ is non-cyclic and requires two generators, $\hat{g}$ and $\hat{h}$. Let $\hat{g}$ be the generator for the positive direction along some axis and $\hat{h}$ the generator for the negative direction. Hence, elements in $G$ are either generated by $\hat{g}$ or $\hat{h}$. This allows us to define the following generating function
$\phi(n)=\begin{cases}\hat{g}^{n},&\text{if }n\geq 0\\ \hat{h}^{n},&\text{otherwise.}\end{cases}$ (18)
This generalizes the orbit definition to
$\operatorname{orb}_{G}(\mathbf{x})$ =$=\{g\mathbf{x}\mid g\in G\}$
=$=\{\phi(n)\mathbf{x}\mid n\in[0,1,\ldots,|G|-1]\}.$ (19)
To further extend this idea, we can also let any function $f(\mathbf{x}):\operatorname{orb}_{G}(\mathbf{x})\to\mathbb{R}$ act on the orbit. We can define $f$ to be an Energy function, $\bar{E}_{\mathbf{\theta}}(\mathbf{x})$. This forms the foundation for our group-induced confidence measure, $\rho_{G}(\mathbf{x},g)$, in Equation 7.
D.2 Proof: Theorem 3.1
Proof.
Let $E_{\mathbf{\theta}}(\mathbf{x})$ be the energy of a $\theta$-parameterised energy model given a point $\mathbf{x}$ as input as defined in Equation 4. Let $\bar{E}_{\mathbf{\theta}}(\mathbf{x})$ be the Monte-Carlo estimate of the energy using Equation 5. As shown in Lemma C.1 we can convolve this energy function on the orbit with a kernel, $K$. To remove visual clutter, we substitute
$\phi(\mathbf{x},g):=\big{(}\bar{E}_{\mathbf{\theta}}\star K\big{)}\big{(}g\mathbf{x}\big{)}.$ (20)
We are interested in the second derivative of $\phi(\mathbf{x},g)$ w.r.t. some $g\in G$. We use the second-order Taylor polynomial and assume sufficiently small remainders. Hence, allowing us to approximate the derivative of $\phi$ by infinitesimal small steps,
$\phi(\mathbf{x},g+h)$ $\approx\phi(\mathbf{x},g)+\phi^{\prime}(\mathbf{x},g)h+\frac{1}{2}\phi^{\prime\prime}(\mathbf{x},g)h^{2}$ (21)
$\phi(\mathbf{x},g-h)$ $\approx\phi(\mathbf{x},g)-\phi^{\prime}(\mathbf{x},g)h+\frac{1}{2}\phi^{\prime\prime}(\mathbf{x},g)h^{2}.$ (22)
Adding both together and rearranging for $\phi^{\prime\prime}$ gives
$\phi(\mathbf{x},g+h)+\phi(\mathbf{x},g-h)\approx 2\phi(\mathbf{x},g)+\phi^{\prime\prime}(\mathbf{x},g)h^{2}$
$\phi^{\prime\prime}(\mathbf{x},g)\approx\frac{1}{h^{2}}\bigg{(}\phi(\mathbf{x},g+h)+\phi(\mathbf{x},g-h)-2\phi(\mathbf{x},g)\bigg{)}.$ (23)
This formulation uses incrementally small steps $g+h$ on the orbit of $G$ and $\mathbf{x}$, which we have to define next.
Given a Matrix Lie Sub-Group, $H$ and a finite sub-group, $G\leq H$. We can approximate $H$ by increasing the cardinality of $G$, such that in the limit $G=H$. Let $\langle g\rangle$ be the generator of $G$, which forms an atomic element of $G$. Then, every $g\in G$ can be expressed by $g=\hat{g}^{k}$ for some $k$. Using composition, a step along the orbit starting at $g$ can be expressed by $(\hat{g}\circ g)\mathbf{x}$. Similarly, a step in the opposite direction can be performed by $(\hat{g}^{-1}\circ g)\mathbf{x}$. We substitute $g\mathbf{x}+h$ by $(\hat{g}\circ g)\mathbf{x}$ in Section D.2.
Let $\mathbf{T}_{g}$ be the matrix representation of $g$. The step length along the orbit is defined by $\|\mathbf{T}_{\hat{g}}-\mathbf{T}_{\hat{g}^{2}}\|$. These two observations lead to the following reformulation of Section D.2,
$\phi^{\prime\prime}(\mathbf{x},g)\approx\frac{1}{\|\mathbf{T}_{\hat{g}}-\mathbf{T}_{\hat{g}^{2}}\|^{2}}$ (24)
$\bigg{(}\phi(\mathbf{x},\hat{g}\circ g)+\phi(\mathbf{x},\hat{g}^{-1}\circ g)-2\phi(\mathbf{x},g)\bigg{)}.$
Remember that $G$ is a finite sub-group of some Matrix Lie Group, $H$. In the limit, however, when the cardinality of $G$ approaches the one of $H$, the approximation error will shrink to zero. Therefore, we consider
$\phi^{\prime\prime}(\mathbf{x},g)=\lim_{|G|\to\infty}\frac{1}{\|\mathbf{T}_{\hat{g}}-\mathbf{T}_{\hat{g}^{2}}\|^{2}}$ (25)
$\bigg{(}\phi(\mathbf{x},\hat{g}\circ g)+\phi(\mathbf{x},\hat{g}^{-1}\circ g)-2\phi(\mathbf{x},g)\bigg{)}.$
This concludes the proof. ∎
D.3 Proof: Theorem 4.1
Proof.
Given a finite sequences $s=g_{0}\circ g_{1}\circ\ldots\circ g_{m}$ of group elements and a sub-sequence $s_{0:k-1}=g_{0}\circ g_{1}\circ\ldots\circ g_{k-1}$ with $k\leq m\in\mathbb{N}$. It is to be shown that additional action by $g_{k}^{-1}$ on $s$ eliminates $g_{k}$ from $s$ (this sequence is coined $s_{\setminus g_{k}}$). We prove the statement $g_{k}^{-1}s=s_{\setminus g_{k}}$ by successive refinement and ignore the trivial case where $k=m$, $k=0$ or $m=0$. We have
$g_{k}^{-1}s=g_{k}^{-1}\circ(g_{0}\circ g_{1}\circ\ldots\circ g_{m}).$ (26)
Due to the associativity axiom of any group, the parentheses can be ignored. Furthermore, $s$ can be substituted by $s_{0:k-1}\circ s_{k:m}$, such that
$g_{k}^{-1}s=g_{k}^{-1}\circ\underbrace{g_{0}\circ\ldots\circ g_{k-1}}_{s_{0:k-1}}\circ\underbrace{g_{k}\circ\ldots\circ g_{m}}_{s_{k:m}}.$ (27)
If elements $s_{0:k-1}$ are non-trivial, then commutativity is required to move $g_{k}^{-1}$ in front of $g_{k}$ to cancel them out. Therefore, we need to assume that $s_{0:k-1}$ and $g_{k}^{-1}$ are elements of an Abelian group. Then,
$g_{k}^{-1}s$ $=s_{0:k-1}\circ g_{k}^{-1}\circ g_{k}\circ\ldots\circ g_{m},$ (28)
$g_{k}^{-1}s$ $=s_{0:k-1}\circ\ldots\circ g_{m},$ (29)
$g_{k}^{-1}s$ $=s_{\setminus g_{k}},$ (30)
which concludes the proof. ∎
D.4 Proof: Markov Process
We begin by formally defining a stochastic and a Markov process.
Definition D.1 (Stochastic Process (Borodin, 2017)).
A stochastic process is a family of random variables $\{X_{t}:t\in\mathcal{T}\}$, where the index $t$ belongs to some arbitrary parameter set, $\mathcal{T}$. Each of the random variables can take the same set of possible values, which are called the state space.
Definition D.2 (Markov Process (Borodin, 2017)).
Given a stochastic process, $\{X_{t}:t\in\mathcal{T}\}$. We say $\{X_{t}:t\in\mathcal{T}\}$ has the Markov property if for all $t$ and all states $\{x_{i}:i\in[0,n]\}$, we have
$\mathbb{P}\left(X_{n+1}=x_{n+1}\mid\{X_{i}=x_{i}:i\in[0,n]\}\right)$
$=\mathbb{P}\left(X_{n+1}=x_{n+1}\mid X_{n}=x_{n}\right).$ (31)
When traversing the search tree, a composition of transformation matrices is formed, $\mathbf{T}_{t}\circ\mathbf{T}_{t+1}\circ\ldots\mathbf{T}_{T}$. These matrices form the state space. At each state, the group-induced confidence measure, $\rho_{G}(\mathbf{x},g)$, is used to determine the next most probable action. To reduce visual clutter, we define $g_{t}:=g\in G_{t}$, i.e., some group element from the group at level $t$. Given a finite set of actions, we use the Boltzmann distribution (softmax) to formulate state-action probabilities,
$\mathbb{P}(\mathbf{T}_{t}\mid\hat{\mathbf{x}})$ $=\frac{\exp\rho_{G_{t}}(\hat{\mathbf{x}},g_{t})}{\sum_{g_{t}}\exp\rho_{G_{t}}(\hat{\mathbf{x}},g_{t})}$ (32)
$=\frac{\exp-\frac{\partial^{2}}{\partial g_{t}^{2}}\big{(}\bar{E}_{\mathbf{\theta}}(g_{t}\hat{\mathbf{x}})\star K\big{)}}{\sum_{g_{t}}\exp-\frac{\partial^{2}}{\partial g_{t}^{2}}\big{(}\bar{E}_{\mathbf{\theta}}(g_{t}\hat{\mathbf{x}})\star K\big{)}},$
where $\hat{\mathbf{x}}$ represents $\mathbf{x}$ after all previous group elements on the trajectory acted on it, i.e., $\hat{\mathbf{x}}:=(g_{t-1}\circ g_{t-2}\circ\ldots\circ g_{0})\mathbf{x}$. The mapping from $\mathbf{x}$ to $\hat{\mathbf{x}}$ is surjective. Trivially, the identity can be injected at all positions without changing $\hat{\mathbf{x}}$. The same is true for (non-trivial) equivariance classes. As long as $\hat{\mathbf{x}}$ stays the same, the conditioning $\mathbb{P}(\mathbf{T}_{g_{t}}\mid\hat{\mathbf{x}})$ will not be affected by the choice of transformation sequence. Due to this finding, we have
$\mathbb{P}(\mathbf{T}_{t}\mid\hat{\mathbf{x}})=\mathbb{P}(\mathbf{T}_{t}\mid\mathbf{T}_{t-1}).$ (33)
This conditioning fulfils the definition of a Markov property.
Next, we complete this proof by showing that we have a stochastic process. $\bar{E}_{\mathbf{\theta}}(\cdot)$ is the expected energy obtained using approximate Bayesian inference,
[…]
where $z$ is a latent scalar for the total energy. As defined in the main corpus of this work, we use Monte-Carlo Dropout to simulate multiple stochastic forward passes, where $\Lambda:=\operatorname{diag}(\boldsymbol{\lambda})\sim\text{Bernoulli}(\lambda\in[0,1])$. So, the measure itself is a random variable sampled from a distribution defined by the neural network. We use kernel density estimation with Gaussian kernels to estimate the energy probabilities, $\mathbb{P}_{\mathbf{\theta}}(z\mid\mathbf{x},\mathbf{\Lambda})$. It is this stochastic measure which renders the process a stochastic process. The transition matrix is formed by $\mathbb{P}(\mathbf{T}_{t}\mid\mathbf{T}_{t-1})$ for all $t$. We are only interested in the forward pass, i.e., traversing the tree from root to leaf. As, however, all group actions are reversible per definition (see Def. 2.1), the backward pass is well-defined, too.